OpenStudy (anonymous):

The sum of two numbers is 20. What is the least possible (minimum) sum of their squares

myininaya (myininaya):

The sum of two numbers is 20 Do you know how to write this as a mathematical equation?

OpenStudy (anonymous):

Im supposed to solve it using completing the squares method

myininaya (myininaya):

Right! So Do you know how to write "the sum of two numbers is 20" as a mathematical equation?

OpenStudy (anonymous):

No, sorry :(

myininaya (myininaya):

Do you know what sum means?

OpenStudy (anonymous):

The answer you get when u add two numbers

myininaya (myininaya):

So "the sum of two numbers is 20" can be written as x+y=20 right?

myininaya (myininaya):

It didn't say what numbers would give us 20 so I just called them x and y

myininaya (myininaya):

so first sentence gives us x+y=20 do you see that?

OpenStudy (anonymous):

Yes, continue

myininaya (myininaya):

What is the least possible (minimum) sum of their squares \[S=x^2+y^2\] So I'm calling the sum of their squares, S

myininaya (myininaya):

Now if x+y=20 then solving this for y we get what?

OpenStudy (anonymous):

Y=20-x .

OpenStudy (anonymous):

x+y=20 x^2+y^2=z z is the min. x^2+y^2=z is the equation of a circle, so the min. is x=y x=10 y=10

myininaya (myininaya):

plug that into our equation we called S

myininaya (myininaya):

\[S=x^2+(20-x)^2\] See I just replace y with 20-x since you said y=20-x

OpenStudy (anonymous):